You are going to be flipping four coins at once: a penny, a nickel, a dime, and a quarter. How many ways are there for the four coins to come up? Also, what is the likelihood of getting four heads on the first toss of each coin?
Answer by stanbon(57214)

You are going to be flipping four coins at once: a penny, a nickel, a dime, and a quarter. How many ways are there for the four coins to come up? Also, what is the likelihood of getting four heads on the first toss of each coin?
Each coin can come up in two ways.
Flipping all four at once allows for 2^4= 16 ways they can come up
Likelyhood of a head on a coin is 1/2.
Likelyhood of 4 heads is (1/2)^4= 1/16
Cheers,
Stan H.